OpenStudy (anonymous):

A quadratic function has its vertex at the point (2,8) The function passes through the point (3,1) Find the quadratic and linear coefficients and the constant term of the function. What formula do I use? I can't remember!

OpenStudy (jtvatsim):

Use, \[y=a(x-h)^2+k\] where (h, k) is the vertex.

OpenStudy (anonymous):

Then how do I find the value for a?

OpenStudy (anonymous):

\[(x-h)^2=2a(y-k)\] \[(x-2)^2=2a(y-8)\] \[x^2-4x+4=2ay-16a\] \[2ay=x^2-4x+4+16a\] \[2ay=x^2-4x+4(1+4a)\] Sub in the the point (3,1) to find a. \[2a=3^2-4(3)+4+4a\] \[2a=-9+12-4\] \[2a=-1\] \[a=-\frac{1}{2}\] Therefore equation is \[(x-2)^2=-(y-8)\]

OpenStudy (anonymous):

Don't know where jtvatsim got that equation from, but mine is the standard form of an equation of a parabola.

OpenStudy (anonymous):

2ay=x2−4x+4(1+4a) In this step, where did you get (1+4a)?

OpenStudy (anonymous):

factorised the 4+16a...

OpenStudy (anonymous):

Which I didn't need to anyway.

OpenStudy (anonymous):

Oh oops. it's 4+16a.

OpenStudy (anonymous):

Let me rewrite the last few lines. Did a boo-boo.

OpenStudy (anonymous):

that's okay! One thing, though, how does one find the linear coefficients and the constant term now?

OpenStudy (anonymous):

\[2a(1)=3^2−4(3)+4+16a\] \[2a=9-12+4+16a\] \[14a=-1\] \[a=-\frac{1}{14}\] therefore \[(x-2)^2=2(-\frac{1}{14})[y-8]\]

OpenStudy (anonymous):

Didn't you read before that I expanded/distributed the brackets in order to find a. You can use that expansion to find your constant. And the "4+16a" is your constant. You just sub in a-value to find your constant.

OpenStudy (anonymous):

\[2ay=x^2−4x+(4+16a)\] \[y=\frac{x^2}{2a}-\frac{2x}{a}+\frac{2(1+4a)}{a}\]

OpenStudy (anonymous):

\[y=(x ^{2}/2a)-(2x/a)+(2(1+4a)/2)\] I'm sorry for seeming like the biggest goof in the world, but I'm stuck here.

OpenStudy (anonymous):

You sub in that a-value I just found for you.

OpenStudy (anonymous):

The last term there is your constant.

OpenStudy (anonymous):

1/2a is your coefficient for x^2. And you can decide for yourself what that coefficient is for the middle term.

OpenStudy (anonymous):

What do you mean I get to decide myself?

OpenStudy (anonymous):

After giving you two answers, you should be capable of knowing the coefficient of x.

OpenStudy (anonymous):

You have been super helpful! It's been a while since I've worked with these types of problems. I'm an English major, so this is not my forte. Thank you!!

OpenStudy (anonymous):

Also, I actually changed the numbers in the problem by the slightest bit, so I wouldn't get the exact answers. I just needed to see the method by which the answers are achieved, not the actual answers. So don't worry, you didn't give them away!

OpenStudy (anonymous):

No worries. Anytime.